### v1

@@ Abstract
Context. Blazars, a class of active galactic nuclei (AGN) powered by supermassive black holes, are known for their remarkable variability across multiple timescales and wavelengths. With advancements in both ground- and space-based telescopes, our understanding of AGN central engines has significantly improved. However, the mechanisms driving this variability remain elusive, and continue to fascinate both theorists and observers alike.
Aims. The primary objective of this study is to constrain the X-ray variability properties of the TeV blazar PKS 2155-304. We conduct a comprehensive X-ray spectral and timing analysis, focusing on both long-term and intra-day variability. This analysis uses data from 22 epochs of XMM-Newton EPIC-pn observations, collected over 15 years (2000 to 2014).
Methods. To investigate the variability of the source, we applied both timing and spectral analyses. For the timing analysis, we estimated fractional variability, variability amplitude, minimum variability timescales, flux distribution, and power spectral density (PSD). In the spectral analysis, we fitted the X-ray spectra using power-law, log-parabola, and broken power-law (BPL) models to determine the best-fitting parameters. Additionally, we studied the hardness ratio (HR).
Results. We observed moderate intra-day variability in most of the light curves. Seven out of the twenty-two observations showed a clear bimodal flux distribution, indicating the presence of two distinct flux states. Our analysis revealed a variable power-law PSD slope. Most HR plots did not show significant variation with flux, except for one observation (OBSID 0124930501), where HR increased with flux (Count/s). The fitted X-ray spectra favored the BPL model for the majority of observations. The findings of this work shed light on the intraday variability of blazars, providing insights into the non-thermal jet processes that drive the observed flux variations.
Key Words.:
@@ 1 Introduction
Blazars belong to a subclass of active galactic nuclei (AGN) characterized by the presence of relativistic jets directed towards our line of sight in a narrow beam of $\leq 10$ degrees (Urry & Padovani, 1995). These sources are distinguished by their high luminosity and flux and polarization variability across diverse time scales. Additionally, they exhibit Doppler-boosted broad-band continuum emission spanning from radio to TeV wavebands. Blazar broadband spectral energy distribution (SED) features two distinct humps in the $\nu$ vs.$\nu f_{\nu}$ plane. The low-energy hump peaks in the optical/X-ray range, while the high-energy hump peaks in the GeV/TeV energies. The lower energy feature is mostly attributed to synchrotron emission from relativistic particles in the magnetic fields present in the jets. However, the origin of the hump at higher energy is still under wide discussion. (see Böttcher, 2019, and references therein). In leptonic models of blazar emission, where the broadband radiative output is considered to be primarily generated by leptons (such as electrons and possibly positrons), the origin of the higher energy feature is most plausibly explained by Inverse-Compton scattering of low-energy seed photons by relativistic particles within the jets. In such a scenario, the seed photons can originate from synchrotron photons within the jets, as proposed in the synchrotron self-Compton model (e.g., Maraschi et al. 1992), and/or from low-energy photons emitted by the accretion disk (Dermer & Schlickeiser, 1993), broad-line region (Sikora et al., 1994), and dusty torus (Błażejowski et al., 2000), commonly referred to as external-Compton (EC) models. For certain blazar sources, the SED suggests a better explanation within hadronic models. In these models, high-energy emission is associated with synchrotron radiation from protons and/or secondary leptons generated through proton-photon interactions (e. g., see Mannheim, 1993; Mücke et al., 2003). One possible way to categorize the blazars is based on whether the rest-frame equivalent width (EW) of their broad optical emission line is less than or greater than 5 Å (Stickel et al., 1991; Stocke et al., 1991). This categorisation introduces two classes of blazars: flat-spectrum radio quasars (FSRQs) and BL Lacertae (BL Lacs). FSRQs are powerful sources dominated by Compton emission, with the synchrotron peak occurring at a lower frequency. In contrast, BL Lac objects, though relatively less powerful, are mostly found to be TeV sources, with their inverse Compton peak located in the highest $\gamma$-ray bands. The dichotomy between BL Lacs and FSRQs is based on intrinsic differences in the nature of the accretion disc and the physical origins of high-energy emissions in these sources. This includes the larger gamma-ray luminosity of FSRQs and the harder gamma-ray spectral characteristics of BL Lacs (see Bhatta et al., 2024, and references therein). Furthermore, in the case of high-energy emission via external inverse-Compton (e. g., Sikora et al., 1994; Dermer & Schlickeiser, 1993), the seed photon field in FSRQs is contributed by different populations of photons external to the jet component, whereas in BL Lacs it is primarily contributed through synchrotron self-Compton (e. g., Ghisellini et al., 1985). Blazars can also be classified based on their location of the synchrotron peak-frequencies ($\nu_{\rm peak}$): low ($\nu_{\rm peak}<10^{14}{\rm\ Hz}$), intermediate ($10^{14}{\rm\ Hz}<\nu_{\rm peak}<10^{15}{\rm\ Hz}$), and high ($\nu_{\rm peak}>10^{15}{\rm\ Hz}$) (Abdo et al., 2010).
@@
Blazars exhibit high-amplitude rapid flux variability across diverse timescales throughout the entire electromagnetic spectrum (see e. g. Bhatta et al., 2023; Bhatta, 2021; Bhatta & Dhital, 2020). This makes multifrequency variability studies particularly effective in exploring the central engine of AGN. The temporal variability can be broadly classified into long-term, short-term, and intraday variability (IDV). Long-term variability, occurring over several months to several years, can lead to flux changes by up to an order of magnitude. Short-term variability, spanning from days to a few months, results in flux changes by a few factors. Additionally, intra-day variability or microvariability are characterized by rapid flux variations within a day (Webb et al., 2021).
@@
In particular, IDV offers insights into the dynamic nature of the relativistic jets and compact regions of blazars, providing strong motivation to explore multi-wavelength IDV in these sources. IDV in several blazars have been studies using the observations in a wide range of energy/frequency bands (see e.g., Aharonian et al., 2007; Bhatta et al., 2013, 2016; Bhatta & Webb, 2018; Aggrawal et al., 2018; Pininti et al., 2023)
@@
Similarly, X-ray variability properties have been extensively studied using observations made by different instruments like XMM- Newton (Dinesh et al., 2023; Mohorian et al., 2022; Noel et al., 2022), Swift (e. g., Kapanadze et al., 2014), Astrosat (Goswami et al., 2024; Markowitz et al., 2022; Bhattacharyya et al., 2020), NuSTAR (e. g., Bhatta et al., 2018), Chandra (e. g., Aggrawal et al., 2018) and RXTE (e. g., Wang et al., 2018). These studies offer crucial insights into blazar variability, including its structure, evolution, dynamics, and the emission mechanisms of the central engine. It is found that the blazar sources exhibit significant flux variability in the X-ray band, with variability timescales ranging from a few minutes to a few hours. In BL Lacs, the observed X-ray emission explores the relativistic population of particles accelerated by turbulence and/or magnetic reconnection prevalent in the jets. In contrast, in FSRQs, the observed emission may significantly originate from thermal instability occurring at the accretion disk. However, the details of the mechanism leading to IDV across diverse EM bands are still under discussion.
@@ 2 Blazar PKS 2155–304
Radio images from the Very Large Array reveal an extended radio jet located approximately 20 kpc from the nucleus of PKS 2155-304 (Liuzzo et al., 2013). The 43 GHz images of the source obtained by very-long-baseline interferometry reveal new morphological details, indicating a significant degree of jet bending within the inner milli-arcsecond of the source (Piner et al., 2010)
The spectral power density of the source’s long-term optical variability is found to be consistent with a broken power-law model, with a characteristic break timescale of approximately 2.7 years (Kastendieck et al., 2011). Moreover, in a study involving long-term optical observations and gamma-ray observations from Fermi/LAT, it was found that the optical emission is highly correlated with the gamma-ray emission (Bhatta, 2021; Rajput et al., 2021). Furthermore, time series analysis of long-term optical and gamma-ray observations of the source revealed quasi-periodic oscillations (QPOs) with characteristic timescales of $\sim$ 600 days and $\sim$ 250 days, which were found to be significant over the red-noise inherent in blazars (Bhatta, 2021).
@@
Multi-wavelength (MWL) analyses of the source have been conducted by modeling the broadband SEDs from different observation epochs using standard blazar models. A log-parabolic SED of PKS 2155–304 was found to be consistent across optical, UV, and X-ray wavelengths (Bhagwan et al., 2014). During a MWL observation campaign spanning from optical to TeV energy ranges, it was discovered that in a low flux state, the object exhibited highly significant flux variability in the X-rays. Additionally, the broadband SED was found to be consistent with a one-zone synchrotron self-Compton model (Abdalla et al., 2020)
The flux and spectral variability of the BL Lac PKS 2155-304 have been extensively studied in the X-ray energy band using observations from several space telescopes, including Chandra, Astrosat, Swift-XRT, NuSTAR, and XMM- Newton (see Gupta, 2020, for a recent review). In a study using NuSTAR observations, Bhatta et al. (2018) reported the presence of steep spectra, with a photon index, $\Gamma\sim 3$, in the hard X-ray emission of the source. A study of the source using X-ray observations from the Suzaku satellite revealed large-amplitude flux and spectral intraday variabilities in highly correlated soft and hard bands, with the blazar exhibiting harder spectra when brighter (Zhang et al., 2021).
@@ 4 ANALYSIS Methods
A comprehensive summary of all the XMM- Newton observations used in our study is presented in Table 1 The table lists the observation ID, instrument, exposure ID, observation mode, observation date, start and end times, and the total duration of each observation. To investigate flux variability, we employed several analytical techniques, each briefly described in this paper. The methodologies are broadly categorized into timing and spectral analyses. The timing analysis involved the estimation of several variability parameters, including excess variance, fractional variability, power spectral density, and flux distribution (see Table 2). For our spectral analysis, we computed hardness ratios by obtaining count rates in two distinct energy bands and calculating their ratios. We then separately fitted the XMM-Newton spectra with three commonly used non-thermal models to find which one is preferred over the 14 years of PKS 2155-304 observations.
@@ 4.1 EXCESS VARIANCE: FLUX VARIABILITY & VARIABILITY AMPLITUDE
where $\bar{\sigma}^{2}_{\rm err}$ represents square of the mean measurement errors. Fractional variability is calculated using the relation,
$F_{\rm var}=\sqrt{\frac{S^{2}-\bar{\sigma}^{2}_{\rm err}}{\bar{X}^{2}}}.$ (2)
The associated error in the fractional variability is obtained using,
$\sigma_{F_{\mathrm{var}}}=\sqrt{\left(\frac{1}{\sqrt{2N}}\frac{\bar{\sigma}^{2}_{err}}{F_{var}}\frac{1}{\bar{X}^{2}}\right)^{2}+\left(\sqrt{\frac{\bar{\sigma}^{2}_{err}}{N}}\frac{1}{\bar{X}^{2}}\right)^{2}}.$ (3)
Variability amplitude $VA$ gives the information about peak to peak flux variations, which can be defined as,
@@
which follows from the propagation of errors in flux in Equation 4. Since in the VA measurement we are only considering peak to peak fluxes, it may not provide the information about the overall variability. In such situation fractional variability, variability index and flux-histograms may provide deeper information about observed flux variability. Since variability of blazars are observed over diverse timescales, we also estimate variability timescale for the PKS 2155–304. The minimum timescale of such variability is (Burbidge et al., 1974; Bhatta et al., 2018),
@@
where $\Delta F_{1}$ and $\Delta F_{2}$ are the flux uncertainties used to estimate the minimum variability timescales for the fluxes $F_{1}$ and $F_{2}$, respectively.
@@ 4.2 RMS–FLUX RELATION
In time-series studies of high-energy astrophysical systems, examining the relationship between the root mean square (RMS) and mean flux offers insights into the nature of the system. In this method of analysis, the source light curve is divided into multiple sections, and the correlation between the RMS and mean flux for each section is analyzed. At first, a linear RMS-flux relationship was observed in the context of variability observed in X-ray binaries hosting accretion discs (e.g., Uttley & McHardy, 2001; Gleissner et al., 2004; Heil et al., 2012), and it was also reported later in the X-ray variability of Seyfert galaxies (Gaskell, 2004; Alston et al., 2019). In recent years, a similar trend has been observed in various frequency bands of blazar observations, including optical, X-ray, and gamma-rays, from several sources (Wang et al., 2023; Kundu et al., 2022; Bhatta, 2021; Bhatta & Dhital, 2020; Bhattacharyya et al., 2020; Edelson et al., 2013)

### v2

@@ Abstract
Blazars, a class of active galactic nuclei (AGN) powered by supermassive black holes, are known for their remarkable variability across multiple timescales and wavelengths. Despite significant advancements in our understanding of AGN central engines, thanks to both ground- and space-based telescopes, the details of the mechanisms driving this variability remain elusive. The primary objective of this study is to constrain the X-ray variability properties of the TeV blazar PKS 2155-304. We conducted a comprehensive X-ray spectral and timing analysis, focusing on both long-term and intra-day variability (IDV), using data from 22 epochs of XMM-Newton observations collected over 15 years (2000 to 2014). For the timing analysis, we estimated the fractional variability, variability amplitude, minimum variability timescales, flux distribution, and power spectral density. In the spectral analysis, we fitted the X-ray spectra using power-law, log-parabola, and broken power-law models to determine the best-fitting parameters. We observed moderate IDV in the majority of the light curves. Seven out of the 22 observations showed a clear bimodal flux distribution, indicating the presence of two distinct flux states. Our analysis revealed a variable power spectral slope. Most hardness ratio plots did not show significant variation with flux, except for two observations, where the hardness ratio changed considerably with flux. The fitted X-ray spectra favored the broken power law model for the majority of observations, indicating break in the spectral profiles. The findings of this work shed light on the IDV of blazars, providing insights into the non-thermal jet processes that drive the observed flux variations.
@@ 1 Introduction
Blazars belong to a subclass of active galactic nuclei (AGN) characterized by the presence of relativistic jets directed towards our line of sight, with an angle between the jet axis and the line of sight of $\leq 10$ degrees (Urry & Padovani, 1995). These sources are distinguished by their high luminosity and flux and polarization variability across diverse timescales. Additionally, they exhibit Doppler-boosted broad-band continuum emission spanning from radio to TeV wavebands. The broadband spectral energy distribution (SED) of blazars features two distinct humps in the $\nu$ vs. $\nu f_{\nu}$ plane. The low-energy hump peaks in the optical/X-ray range, whereas the high-energy hump peaks in the GeV/TeV energy range. The lower energy feature is mostly attributed to synchrotron emission from relativistic particles in the magnetic fields present in the jets. However, the origin of the hump at higher energy is still under wide discussion (see Böttcher, 2019, and references therein). In leptonic models of blazar emission, where the broadband radiative output is considered to be primarily generated by leptons (such as electrons and possibly positrons), the origin of the high-energy feature is most plausibly explained by Inverse-Compton scattering of low-energy seed photons by relativistic particles within the jets. In such a scenario, the seed photons can originate from synchrotron photons within the jets, as proposed in the synchrotron self-Compton model (e.g., Maraschi et al. 1992), and/or from low-energy photons emitted by the accretion disk (Dermer & Schlickeiser, 1993), broad-line region (Sikora et al., 1994), and dusty torus (Błażejowski et al., 2000), collectively referred to as external-Compton (EC) models. For certain blazar sources, the SED suggests a better explanation within hadronic models. In these models, high-energy emission is associated with synchrotron radiation from protons and/or secondary leptons generated through proton-photon interactions (e. g., see Mannheim, 1993; Mücke et al., 2003). One possible way to categorize the blazars is based on whether the rest-frame equivalent width (EW) of their broad optical emission line is less than or greater than 5 Å (Stickel et al., 1991; Stocke et al., 1991). This categorization introduces two classes of blazars: flat-spectrum radio quasars (FSRQs) and BL Lacertae (BL Lacs).
FSRQs are powerful sources dominated by Compton emission, with the synchrotron peak occurring at a lower frequency. In contrast, BL Lac objects, though relatively less powerful, are mostly found to be TeV sources, with their inverse Compton peak located in the highest $\gamma$-ray bands. The dichotomy between BL Lacs and FSRQs is based on intrinsic differences in the nature of the accretion disc and the physical origins of high-energy emissions in these sources. This includes the larger gamma-ray luminosity of FSRQs and the harder gamma-ray spectral characteristics of BL Lacs (see Bhatta et al., 2024, and references therein). Furthermore, in the case of high-energy emission via external inverse-Compton (e. g., Sikora et al., 1994; Dermer & Schlickeiser, 1993), the seed photon field in FSRQs is contributed by different populations of photons external to the jet component, whereas in BL Lacs it is primarily contributed through synchrotron self-Compton (e. g., Ghisellini et al., 1985). Blazars can also be classified based on their location of the synchrotron peak-frequencies ($\nu_{\rm peak}$): low ($\nu_{\rm peak}<10^{14}{\rm\ Hz}$), intermediate ($10^{14}{\rm\ Hz}<\nu_{\rm peak}<10^{15}{\rm\ Hz}$), and high ($\nu_{\rm peak}>10^{15}{\rm\ Hz}$) (Abdo et al., 2010). Moreover, high-synchrotron peaked blazars with synchrotron peak frequencies above 1017 Hz are specifically classified as extreme high-energy synchrotron peaked blazars (see Chang et al., 2019, and references therein).
@@
Blazars exhibit high-amplitude rapid flux variability across diverse timescales throughout the entire electromagnetic spectrum (see e. g, Bhatta et al., 2023; Bhatta, 2021; Bhatta & Dhital, 2020). This makes multifrequency variability studies particularly effective in exploring the central engine of AGN. The temporal variability can be broadly classified into long-term, short-term, and intra-day variability (IDV). Long-term variability, occurring over several months to several years, can lead to flux changes by up to an order of magnitude (e. g., Kastendieck et al., 2011; Rajput et al., 2020; Bhatta, 2021). Short-term variability, spanning days to a few months, results in flux changes by a few factors (e. g., Pininti et al., 2023; Bhatta et al., 2023; Foschini et al., 2006). Furthermore, IDV or microvariability is characterized by rapid flux variations within a day (Webb et al., 2021).
@@
In particular, IDV offers insights into the dynamic nature of the relativistic jets and compact regions of blazars, providing strong motivation to explore multi-wavelength IDV in these sources. IDV in several blazars has been studied using observations in a wide range of energy/frequency bands (see e.g., Aharonian et al., 2007; Bhatta et al., 2013, 2016; Bhatta & Webb, 2018; Aggrawal et al., 2018; Pininti et al., 2023).
@@
Similarly, X-ray variability properties have been extensively studied using observations made by different instruments like XMM-Newton (Dinesh et al., 2023; Mohorian et al., 2022; Noel et al., 2022), Swift (e. g., Kapanadze et al., 2014), Astrosat (Goswami et al., 2024; Markowitz et al., 2022; Bhattacharyya et al., 2020), NuSTAR (e. g., Bhatta et al., 2018), Chandra (e. g., Aggrawal et al., 2018) and RXTE (e. g., Wang et al., 2018). These studies offer crucial insights into blazar variability, including its structure, evolution, dynamics, and the emission mechanisms of the central engine. Blazar sources are found to exhibit significant flux variability in the X-ray band, with variability timescales ranging from a few minutes to a few decades. In BL Lac objects, the observed X-ray emission probes the relativistic particle population accelerated by internal shocks (e. g., Böttcher & Dermer, 2010), turbulence, and/or magnetic reconnection (e. g., Marscher & Jorstad, 2021; Christie et al., 2019) prevalent within their jets. In contrast, FSRQs exhibit substantial thermal emission from their accretion disks, which predominantly appears as the characteristic ‘big blue bump’ in the optical/UV bands (e. g., Paltani et al., 1998). As a result, the observed X-ray emission may partly arise from thermal instabilities within the accretion disk structures (see Jolley et al., 2009, for discussion on disk-jet connection in blazars). However, the details of the mechanism leading to IDV across diverse EM bands are still under discussion.
As part of a broader program investigating the MWL emission mechanisms in blazars, this work focuses on characterizing the X-ray variability of PKS 2155-304, a well-studied HSP blazar, through detailed multi-timescale analysis. Our primary objective is to characterize both the intra-day and long-term variability properties of this source through detailed X-ray observations (e. g., see our similar previous works on other sources Mohorian et al., 2022; Bhatta et al., 2018; Dinesh et al., 2023). By examining the temporal and spectral behavior across different timescales, we aim to better understand the physical mechanisms driving the MWL variability and high-energy emission in blazars. Blazar PKS 2155-304 serves as an excellent laboratory for this investigation due to its strong X-ray emission and documented variability across the electromagnetic spectrum, making it an ideal candidate for probing the underlying physics of blazar emission mechanisms and jet dynamics.
@@ 2 Blazar PKS 2155–304
Radio images from the Very Large Array reveal an extended radio jet located approximately 20 kpc from the nucleus of PKS 2155-304 (Liuzzo et al., 2013). The 43 GHz images of the source obtained by very-long-baseline interferometry reveal new morphological details, indicating a significant degree of jet bending within the inner milli-arcsecond of the source (Piner et al., 2010).
The spectral power density of the source’s long-term optical variability is found to be consistent with a broken power-law model, with a characteristic break timescale of approximately 2.7 years (Kastendieck et al., 2011). Moreover, in a study involving long-term optical observations and gamma-ray observations from Fermi/LAT, it was found that the optical emission is highly correlated with the gamma-ray emission (Bhatta, 2021; Rajput et al., 2021). Furthermore, time series analysis of long-term optical and gamma-ray observations of the source revealed quasi-periodic oscillations (QPOs) with characteristic timescales of $\sim$ 600 days and $\sim$ 250 days, which were found to be significant over the red-noise inherent in blazars (Bhatta, 2021; Peñil et al., 2020; Chevalier et al., 2019; Sandrinelli et al., 2014).
@@
Multiwavelength (MWL) analyses of the source have been conducted by modeling the broadband SEDs from different observation epochs using standard blazar models. A log-parabolic SED of PKS 2155–304 was found to be consistent across optical, UV, and X-ray wavelengths (Bhagwan et al., 2014). During a MWL observation campaign spanning from optical to TeV energy ranges, it was discovered that in a low flux state, the object exhibited highly significant flux variability in the X-rays. Additionally, the broadband SED was found to be consistent with a one-zone synchrotron self-Compton model (Abdalla et al., 2020).
The flux and spectral variability of the BL Lac PKS 2155-304 have been extensively studied in the X-ray energy band using observations from several space telescopes, including Chandra, Astrosat, Swift-XRT, NuSTAR, and XMM-Newton (see Gupta, 2020, for a recent review). In a study using NuSTAR observations, Bhatta et al. (2018) reported the presence of steep spectra, with a photon index, $\Gamma\sim 3$, in the hard X-ray emission of the source. A study of the source using X-ray observations from the Suzaku satellite revealed large-amplitude flux and spectral intra-day variabilities in highly correlated soft and hard bands, with the blazar exhibiting harder spectra when brighter (Zhang et al., 2021).
@@ 4 ANALYSIS Methods
A comprehensive summary of all the XMM-Newton observations used in our study is presented in Table 1. The table lists the observation ID, instrument, exposure ID, observation mode, observation date, start and end times, and the total duration of each observation. To investigate flux variability, we employed several analytical techniques, each briefly described in this paper. The methodologies are broadly categorized into timing and spectral analyses. The timing analysis involved the estimation of several variability parameters, including excess variance, fractional variability, power spectral density, and flux distribution (see Table 2). For our spectral analysis, we computed hardness ratios by obtaining count rates in two distinct energy bands and calculating their ratios. We then separately fitted the XMM-Newton spectra with three commonly used non-thermal models to find which one is preferred over the 14 years of PKS 2155-304 observations.
@@ 4.1 EXCESS VARIANCE: FLUX VARIABILITY & VARIABILITY AMPLITUDE
where $\bar{\sigma}^{2}_{\rm err}$ represents square of the mean measurement errors (Nandra et al., 1997; Edelson et al., 2002). Fractional variability is calculated using the relation,
$F_{\rm var}=\sqrt{\frac{S^{2}-\bar{\sigma}^{2}_{\rm err}}{\bar{X}^{2}}}$ (2)
(Edelson et al., 1990; Vaughan et al., 2003). The associated error in the fractional variability is obtained using,
$\sigma_{F_{\mathrm{var}}}=\sqrt{\left(\frac{1}{\sqrt{2N}}\frac{\bar{\sigma}^{2}_{err}}{F_{var}}\frac{1}{\bar{X}^{2}}\right)^{2}+\left(\sqrt{\frac{\bar{\sigma}^{2}_{err}}{N}}\frac{1}{\bar{X}^{2}}\right)^{2}}$ (3)
(Aleksić et al., 2015; Bhatta & Webb, 2018). Variability amplitude $VA$ gives the information about peak-to-peak flux variations, which can be defined as,
@@
which follows from the propagation of errors in flux in Equation 4 (Bhatta et al., 2018; Mohorian et al., 2022). Since in the VA measurement we are only considering peak to peak fluxes, it may not provide the information about the overall variability. In such situation fractional variability, variability index and flux-histograms may provide deeper information about observed flux variability. Since variability of blazars are observed over diverse timescales, we also estimate variability timescale for the PKS 2155–304. The minimum timescale of such variability is (Burbidge et al., 1974)
@@
where $\Delta F_{1}$ and $\Delta F_{2}$ are the flux uncertainties used to estimate the minimum variability timescales for the fluxes $F_{1}$ and $F_{2}$, respectively (see Bhatta et al., 2018).
@@ 4.2 RMS–FLUX RELATION
In time-series studies of high-energy astrophysical systems, examining the relationship between the root mean square (RMS) and mean flux offers insights into the nature of the system. In this method of analysis, the source light curve is divided into multiple sections, and the correlation between the RMS and mean flux for each section is analyzed. At first, a linear RMS-flux relationship was observed in the context of variability observed in X-ray binaries hosting accretion discs (e.g., Uttley & McHardy, 2001; Gleissner et al., 2004; Heil et al., 2012), and it was also reported later in the X-ray variability of Seyfert galaxies (Gaskell, 2004; Alston et al., 2019). In recent years, a similar trend has been observed in various frequency bands of blazar observations, including optical, X-ray, and gamma-rays, from several sources (Wang et al., 2023; Kundu et al., 2022; Bhatta, 2021; Bhatta & Dhital, 2020; Bhattacharyya et al., 2020; Edelson et al., 2013).